# Planet formation theory: an overview

### v1

@@ 6 Stability and planetary system evolution
  • Gas in the co-orbital region exchanges angular momentum with the planet as it executes horseshoe turns in the frame co-rotating with the planet. Asymmetries in the behavior of the turns in the regions that lead / trail the planet, due to radial gradients in entropy and vortensity, lead to a net torque. Complexity arise because, in the absence of viscosity or wind-driven radial motion, the streamlines in the horseshoe region are closed, and would rapidly adjust to null out (“saturate”) the torque. Simulations show that adding the co-orbital torque to the Lindblad torque can lead to outward migration, but that this result depends not just on the disk radial structure, but also on how energy and angular momentum are transported (Paardekooper et al. 2010, 2011).
@@
In low-viscosity disks, the picture is more complex. Planet-disk interaction in this regime commonly forms vortices through the Rossby Wave Instability (Lovelace et al. 1999), which can themselves migrate and interact with the planet. McNally et al. (2019) show how the migration regime varies in a two-dimensional plane, with the planet mass and disk viscosity as the controlling parameters. Particularly interesting is the case of a disk with a low turbulent viscosity, but efficient wind angular momentum loss. Wind angular momentum loss is essentially a surface rather than a volume process, so a wind-driven disk is expected to drive a rapid radial flow through planetary gaps that will affect migration torques (McNally et al. 2020; Wu and Chen 2025; Hu et al. 2024). Using three dimensional non-ideal magnetohydrodynamic simulations, Wafflard-Fernandez and Lesur (2023) find that magnetic flux concentration in the gap region leads to an asymmetric gap profile that leads to outward migration of massive planets.
@@
In the linear, Type I regime, the disk torque on low mass planets is expected to be largely independent of the multiplicity of the planetary system. The migration of the system will, however, be affected if there is angular momentum exchange due to mean-motion resonances between the planets. Resonant capture (Goldreich and Peale 1966; Borderies and Goldreich 1984) occurs when there is convergent migration into a mean motion resonance, with a probability that depends on the migration speed, eccentricity, and strength of perturbations (due, in the case of migrating planets, to gravitational coupling to disk turbulence) (Quillen 2006; Adams et al. 2008; Batygin 2015). For giant planets in the Type II regime, overlap of the gaps created by the planets means that the Lindblad torques acting on them are no longer independent. Under the right conditions, two giant planets, that individually would each migrate inward, migrate outward as a coupled resonant pair (Masset and Snellgrove 2001). These rather subtle multi-planet effects are of central importance to the “breaking the chains” model for close-in Kepler systems (Izidoro et al. 2017), for modern version of the Nice model for the outer Solar System (Tsiganis et al. 2005; Morbidelli et al. 2007), and in the Grand Tack model for the inner Solar System (Walsh et al. 2011). Planetesimal scattering, subsequent to gas disk dispersal, provides a mechanism for breaking resonances, especially in the context of the Nice model. Given a sufficiently massive planetesimal disk, scattering may be important across a broader range of radii during planet formation (Minton and Levison 2014; Jinno et al. 2024).
@@
Migration is a phenomenon that, once protoplanets exceed masses of a fraction of an Earth mass, impacts all phases of planet formation and planetary system evolution. It has a similar relation to planet formation as opacity does to stellar structure, in that it is critically important, complex, but in principle calculable from first principles to useful precision. There are only a few circumstances where completely ignoring migration is a defensible position.
@@
After the dissipation of the protoplanetary disk, planets with atmospheres or envelopes can lose mass through photoevaporation (Owen and Wu 2013) and / or core-powered mass loss (Ginzburg et al. 2016). The immediate architecture emerging from the disk phase – during which planet-disk interactions typically (though not always) damp eccentricity and inclination – may also prove to be unstable over long time scales. The simplest model problem here is the evolution of a system of $N$ planets, initially on circular, coplanar orbits, under Newtonian point-mass gravity. For $N=2$, stability to close encounters between the planets (“Hill stability”) can be proven beyond a threshold separation (Gladman 1993). For $N\geq 3$, planetary systems are found through N-body integrations to be unstable, with a time scale to instability that is a steep function of the initial separation (measured in units of the planet mass to the 1/3 or 1/4 power; Chambers et al. 1996; Smith and Lissauer 2009; Obertas et al. 2017). These numerical findings are interpreted within a dynamical framework of chaos driven by the overlap of resonances (Lecar et al. 2001; Quillen 2011; Lammers et al. 2024). Resonant chain systems – arguably the most natural initial condition emerging from the disk phase – can of course have much longer lifetimes than similar planetary systems with non-resonant period ratios. Pichierri and Morbidelli (2020) and Goldberg et al. (2022) analyze the stability of this important special case.
@@
The outcome of dynamical instability in a planetary system can be ejection of one or more of the planets, physical collisions, or excitation to a high eccentricity state that may collide or interact tidally with the host star. The parameter that controls the branching ratios between these outcomes is the ratio of the orbital velocity to the escape speed from the surface of the planet. High values of $v_{\rm K}/v_{\rm esc}$ (low mass planets, and / or small orbital radii) favor collisions, while low values lead to a predominance of ejections (Ford et al. 2001; Petrovich et al. 2014).
@@
where the sum is over planets with angular momenta $\Lambda_{k}$, semi-major axes $a_{k}$, eccentricities $e_{k}$, and inclinations (relative to the invariable plane) $i_{k}$. AMD conservation ensures that nothing interesting develops from initially near-circular, near-coplanar systems of widely separated planets (“AMD stability”; Laskar and Petit 2017). Given non-zero AMD, however (potentially arising from prior planet-planet scattering events), chaos driven by secular resonances leads to each planet eventually experiencing substantial excursions in AMD about an equipartition value (Wu and Lithwick 2011). A large AMD for an inner, relatively low mass planet, can lead to a high eccentricity that places the pericenter distance close enough to the star for tidal dissipation to become important. Qualitatively similar behavior is possible under Kozai-Lidov dynamics (Kozai 1962; Lidov 1962; Naoz et al. 2011), in which a planet exhibits cyclic variations in eccentricity due to secular perturbations from a distant, inclined companion. These secular processes can naturally produce high (and even retrograde) inclinations as well as high eccentricities (Wu and Murray 2003; Fabrycky and Tremaine 2007; Naoz et al. 2011), making them prime suspects in the formation of the misaligned exoplanets that have been detected via the Rossiter-Mclaughlin effect (Winn et al. 2010).
@@ 7 Some open questions
    Uncertainty quantification of migration rates. Computing the rate of gas disk migration to a reliably known level of precision, across the range of planet masses and disk structures relevant to planet formation, would remove a major source of uncertainty. Code comparison efforts are a valuable step toward this goal.
@@
    Is planet formation smoothly distributed or radially concentrated at special disk locations? Protoplanetary disks contain ice lines, and probably other deterministic radial locations where the disk physics creates particle traps or planet migration null points. The basic paradigm for planet formation could yet be either smooth or ring-like.
@@
    Do disks evolve mostly due to turbulent or wind torques? Magnetized winds are currently favored, but even if the basic question is considered resolved, the strength and nature of the turbulence in primarily wind-driven disks is still of fundamental importance for planet formation.
@@
    Does the composition of the upper levels of exoplanet atmospheres retain formation information? The division of abundant elements such as carbon and oxygen between condensed and gaseous phases is expected to vary with temperature, and thus radius, in the disk. This opens up the possibility of using atmospheric composition measurements from transmission spectroscopy as a new probe of planet formation (Öberg et al. 2011). Multiple confounding factors, in both planet formation and atmospheric evolution, might however frustrate such hopes.
@@
    How can theoretical models of planetary system evolution be tested against population-level data? Even when the theoretical model is well-specified and accurately calculable (e.g. pure N-body dynamics), and the data well-understood, inferring principled statistical constraints is often difficult. New statistical and Machine Learning approaches show promise.
@@ Acknowledgements.
My understanding of planet formation has been greatly informed by the work and ideas of students, postdocs, and other collaborators over many years. Thanks to all of you!

### v2

@@ 6 Stability and planetary system evolution
  • Gas in the co-orbital region exchanges angular momentum with the planet as it executes horseshoe turns in the frame co-rotating with the planet. Asymmetries in the behavior of the turns in the regions that lead / trail the planet, due to radial gradients in entropy and vortensity, lead to a net torque. Complexity arises because, in the absence of viscosity or wind-driven radial motion, the streamlines in the horseshoe region are closed, and would rapidly adjust to null out (“saturate”) the torque. Simulations show that adding the co-orbital torque to the Lindblad torque can lead to outward migration, but that this result depends not just on the disk radial structure, but also on how energy and angular momentum are transported (Paardekooper et al. 2010, 2011).
@@
In low-viscosity disks, the picture is more complex. Planet-disk interaction in this regime commonly forms vortices through the Rossby Wave Instability (Lovelace et al. 1999), which can themselves migrate and interact with the planet. McNally et al. (2019) show how the migration regime varies in a two-dimensional plane, with the planet mass and disk viscosity as the controlling parameters. Particularly interesting is the case of a disk with a low turbulent viscosity, but efficient wind angular momentum loss. Wind angular momentum loss is essentially a surface rather than a volume process, so a wind-driven disk is expected to drive a rapid radial flow through planetary gaps that will affect migration torques (McNally et al. 2020; Kimmig et al. 2020; Speedie et al. 2022; Wu and Chen 2025; Hu et al. 2024). Using three dimensional non-ideal magnetohydrodynamic simulations, Wafflard-Fernandez and Lesur (2023) find that magnetic flux concentration in the gap region leads to an asymmetric gap profile that leads to outward migration of massive planets.
@@
In the linear, Type I regime, the disk torque on low mass planets is expected to be largely independent of the multiplicity of a planetary system. The migration of individual planets in a system will, however, be affected if there is angular momentum exchange due to mean-motion resonances between the planets. Resonant capture (Goldreich and Peale 1966; Borderies and Goldreich 1984) occurs when there is convergent migration into a mean motion resonance, with a capture probability that depends on the migration speed, eccentricity, and strength of perturbations (due, in the case of migrating planets, to gravitational coupling to disk turbulence) (Quillen 2006; Adams et al. 2008; Batygin 2015). For giant planets in the Type II regime, an additional effect arises because overlap of the gaps created by the planets means that the Lindblad torques acting on them are no longer independent. Under the right conditions, two giant planets, that individually would each migrate inward, migrate outward as a coupled resonant pair (Masset and Snellgrove 2001). These rather subtle multi-planet effects are of central importance to the “breaking the chains” model for close-in Kepler systems (Izidoro et al. 2017), for modern version of the Nice model for the outer Solar System (Tsiganis et al. 2005; Morbidelli et al. 2007), and in the Grand Tack model for the inner Solar System (Walsh et al. 2011). Planetesimal scattering, subsequent to gas disk dispersal, provides a mechanism for breaking resonances, especially in the context of the Nice model. Given a sufficiently massive planetesimal disk, scattering may be important across a broader range of radii during planet formation (Minton and Levison 2014; Jinno et al. 2024).
@@
Migration is a phenomenon that, once protoplanets exceed masses of a fraction of an Earth mass, impacts all phases of planet formation and planetary system evolution. It has a similar relation to planet formation as opacity does to stellar structure, in that it is critically important, complex, but in principle calculable from first principles to useful precision. There are only a few circumstances where completely ignoring migration in planet formation is a defensible position.
@@
After the dissipation of the protoplanetary disk, planets with atmospheres or envelopes can lose mass through photoevaporation (Owen and Wu 2013) and / or core-powered mass loss (Ginzburg et al. 2016). The immediate architecture emerging from the disk phase – during which planet-disk interactions typically (though not always) damp eccentricity and inclination – may also prove to be unstable over long time scales. The simplest model problem here is the evolution of a system of $N$ planets, initially on circular, coplanar orbits, under Newtonian point-mass gravity. For $N=2$, stability to close encounters between the planets (“Hill stability”) can be proven beyond a threshold separation (Gladman 1993). For $N\geq 3$, model planetary systems with equal mass planets, uniformly spaced in $\log(a)$, are found through N-body integrations to be unstable, with a time scale to instability that is a steep function of the initial separation (Chambers et al. 1996; Smith and Lissauer 2009; Obertas et al. 2017). Although a variety of empirical functions have been proposed, recent results suggest that the instability time is best fit by a power-law function of separation, where the separation is measured in units of the planet mass to the 1/4 power (Lammers et al. 2024). These numerical findings are interpreted within a dynamical framework of chaos driven by the overlap of resonances (Lecar et al. 2001; Quillen 2011; Yalinewich and Petrovich 2020; Petit et al. 2020; Lammers et al. 2024). Resonant chain systems – arguably the most natural initial condition emerging from the disk phase – have much longer lifetimes than similar planetary systems with non-resonant period ratios. Pichierri and Morbidelli (2020) and Goldberg et al. (2022) analyze the stability of this important special case.
@@
The outcome of dynamical instability in a planetary system can be ejection of one or more of the planets, physical collisions, or excitation to a high eccentricity state that may collide or interact tidally with the host star. The parameter that controls the branching ratios between these outcomes is the ratio of the orbital velocity to the escape speed from the surface of the planet. High values of $v_{\rm K}/v_{\rm esc}$ (low mass planets, and / or small orbital radii) favor collisions, while low values lead to a predominance of ejections (Ford et al. 2001; Petrovich et al. 2014). Collisions dominate for systems of close-in super-Earth and sub-Neptune mass planets, whose properties resemble the outcomes seen in simulations of richer, initially unstable, planetary systems (Pu and Wu 2015).
@@
where the sum is over planets with angular momenta $\Lambda_{k}$, semi-major axes $a_{k}$, eccentricities $e_{k}$, and inclinations (relative to the invariable plane) $i_{k}$. AMD conservation ensures that no interesting dynamics develops from initially near-circular, near-coplanar systems of widely separated planets (“AMD stability”; Laskar and Petit 2017). Given non-zero AMD, however (potentially arising from prior planet-planet scattering events), chaos driven by secular resonances leads to each planet eventually experiencing substantial excursions in AMD about an equipartition value (Wu and Lithwick 2011). A large AMD for an inner, relatively low mass planet, can lead to a high eccentricity that places the pericenter distance close enough to the star for tidal dissipation to become important. Qualitatively similar behavior is possible under Kozai-Lidov dynamics (Kozai 1962; Lidov 1962; Naoz et al. 2011), in which a planet exhibits cyclic variations in eccentricity due to secular perturbations from a distant, inclined companion. These secular processes can naturally produce high (and even retrograde) inclinations as well as high eccentricities (Wu and Murray 2003; Fabrycky and Tremaine 2007; Naoz et al. 2011), making them prime suspects in the formation of the misaligned exoplanets that have been detected via the Rossiter-Mclaughlin effect (Winn et al. 2010).
@@ 7 Some open questions
    Uncertainty quantification of planet migration rates. Computing the rate of gas disk migration to a reliably known level of precision, across the range of planet masses and disk structures relevant to planet formation, would remove a major source of uncertainty in planet formation theory. Code comparison efforts are a valuable step toward this goal.
@@
    Is planet formation smoothly distributed or radially concentrated at special disk locations? Protoplanetary disks contain ice lines, and probably other deterministic radial locations where the disk physics creates particle traps or planet migration null points. How important these “special” locations are remains rather unclear. Put simply, the zeroth order paradigm for planet formation could yet be either smooth or ring-like.
@@
    Do disks evolve mostly due to turbulent or wind torques? Magnetized winds are currently favored, but even if the basic question is considered resolved, the strength and nature of the turbulence in primarily wind-driven disks is still of fundamental importance. It matters, among other things, for planetesimal formation, pebble accretion, and planet migration.
@@
    Does the composition of the upper levels of exoplanet atmospheres retain formation information? The division of abundant elements such as carbon and oxygen between condensed and gaseous phases is expected to vary with temperature, and thus radius, in the disk. This opens up the possibility of using atmospheric composition measurements from transmission spectroscopy as a new probe of planet formation (Öberg et al. 2011). Multiple confounding factors, in both planet formation and atmospheric evolution, might however frustrate such hopes. The Chapter by Pudritz, Cridland, and Inglis reviews this question in depth.
@@
    How can theoretical models of planetary system evolution be tested against population-level data? At the most ambitious level, this requires population synthesis, reviewed in this volume by Burn and Mordasini, in which all sources of theoretical and observational uncertainty must be considered. However, even in more limited comparisons when the theoretical model is well-specified and accurately calculable (e.g. pure N-body dynamics), and the data well-understood, inferring principled statistical constraints is often difficult. New statistical and Machine Learning approaches show promise.
@@
    What combination of processes lead to planetesimal formation? Is the streaming instability able to form planetesimals – across the range of disk radii and time necessary for planet formation – on its own, or are other processes (such as particle concentration in large-scale vortices) that “pre-concentrate” solids necessary?
  10. 10.
@@ Acknowledgements.
My understanding of planet formation has been greatly informed by the work and ideas of students, postdocs, and other collaborators over many years. Thanks to all of you! My work has been supported by NASA, the NSF, and the Simons Foundation. I thank Ralph Pudritz for a careful review of this Chapter.